# Optimal Control of a Subdiffusion model using Dirichlet-Neumann and Neumann-Neumann Waveform Relaxation Algorithms

### v1

@@ 4 Auxiliary Results
$=\begin{bmatrix}I_{m,m}&O_{m,m}\\ O_{m,m}&I_{m,m}\end{bmatrix}+\frac{1}{2}\begin{bmatrix}-L_{m,m}-J_{m,m}L_{m,m}J_{m,m}+J_{m,m}B_{m,m}&-L_{m,m}J_{m,m}+J_{m,m}L_{m,m}+J_{m,m}B_{m,m}J_{m,m}\\ -J_{m,m}L_{m,m}+L_{m,m}J_{m,m}-B_{m,m}&-L_{m,m}-J_{m,m}L_{m,m}J_{m,m}-B_{m,m}J_{m,m}\end{bmatrix}.$
Now the matrix block takes the form:
$-L_{m,m}-J_{m,m}L_{m,m}J_{m,m}+J_{m,m}B_{m,m}$ $=\begin{bmatrix}&-1&&&\\ -1&&-1&&\\ &\ddots&&\ddots&\\ &&-1&&-1\\ &&&-1&-1\end{bmatrix},$
$-L_{m,m}-J_{m,m}L_{m,m}J_{m,m}-B_{m,m}J_{m,m}$ $=\begin{bmatrix}1&-1&&&\\ -1&&-1&&\\ &\ddots&&\ddots&\\ &&-1&&-1\\ &&&-1&\end{bmatrix},$
$-L_{m,m}J_{m,m}+J_{m,m}L_{m,m}+J_{m,m}B_{m,m}J_{m,m}$ $=\begin{bmatrix}&&&1&\\ &&1&&-1\\ &\iddots&&\iddots&\\ 1&&-1&&\\ -1&-1&&&\end{bmatrix},$
$-J_{m,m}L_{m,m}+L_{m,m}J_{m,m}-B_{m,m}$ $=\begin{bmatrix}&&&-1&1\\ &&-1&&1\\ &\iddots&&\iddots&\\ -1&&1&&\\ &1&&&\end{bmatrix}.$
[…]
$X^{T}SX$
$=\begin{bmatrix}I_{m,m}&O_{m,1}&O_{m,m}\\ O_{1,m}&I_{1.1}&O_{1,m}\\ O_{m,m}&O_{m,1}&I_{m,m}\end{bmatrix}+\frac{1}{2}\begin{bmatrix}-L_{m,m}-J_{m,m}L_{m,m}J_{m,m}&\sqrt{2}J_{m,m}D_{m,1}&-L_{m,m}J_{m,m}+J_{m,m}L_{m,m}\\ \sqrt{2}B_{1,m}&O_{1,1}&\sqrt{2}B_{1,m}J_{m,m}\\ -J_{m,m}L_{m,m}+L_{m,m}J_{m,m}&\sqrt{2}D_{m,1}&-J_{m,m}L_{m,m}J_{m,m}-L_{m,m}\end{bmatrix}.$
Now using the similar arguments as in the even case we can conclude that each Greshgorin radius $R_{i}\leq 2$.
Corollary 4.19.
Let $S$ be the Caputo finite difference matrix as in corollary 4.13 and $n$ be even. Then
$C_{m,m}$ $=\begin{bmatrix}d_{m+1,1}&&&&\\ d_{m+2,2}-d_{m+2,1}&d_{m+2,1}&&&\\ \vdots&&\ddots&&\\ d_{n-1,m-1}-d_{n-1,m-2}&\cdots&&d_{n-1,1}&\\ d_{n,m}-d_{n,m-1}&\cdots&&d_{n,2}-d_{n,1}&d_{n,1}\end{bmatrix},$
$B_{m,m}$ $=\begin{bmatrix}d_{m+1,m+1}-d_{m+1,m}&&\cdots&&d_{m+1,2}-d_{m+1,1}\\ \vdots&&\vdots&&\vdots\\ d_{n,n}-d_{n,n-1}&&\cdots&&d_{n,m+1}-d_{n,m}\end{bmatrix}.$
Therefore from the relation (14) different blocks of $X^{T}SX$ matrix are:
$A_{m,m}+J_{m,m}C_{m,m}J_{m,m}+J_{m,m}B_{m,m}$
$=\begin{bmatrix}d_{1,1}+d_{n,1}&d_{n,2}-d_{n,1}&&\cdots&d_{n,m}-d_{n,m-1}\\ d_{2,2}-d_{2,1}&d_{2,1}+d_{n-1,1}&&\cdots&d_{n-1,m-1}-d_{n-1,m-2}\\ \vdots&&\ddots&&\vdots\\ d_{m-1,m-1}-d_{m-1,m-2}&&\cdots&d_{m-1,1}+d_{m+2,1}&d_{m+2,2}-d_{m+2,1}\\ d_{m,m}-d_{m,m-1}&&\cdots&d_{m,2}-d_{m,1}&d_{m,1}+d_{m+1,1}\end{bmatrix}$
$+\begin{bmatrix}d_{n,n}-d_{n,n-1}&&\cdots&&d_{n,m+1}-d_{n,m}\\ \vdots&&\vdots&&\vdots\\ d_{m+1,m+1}-d_{m+1,m}&&\cdots&&d_{m+1,2}-d_{m+1,1}\end{bmatrix},$
For computational simplicity we take the Gershgorin radius from this block for the center $\underline{C}_{1}=d_{1,1}+d_{n,1},\underline{C}_{2}=d_{2,1}+d_{n-1,1},\cdots,\underline{C}_{m}=d_{m,1}+d_{m+1,1}$ are
$\underline{R}_{1}$ $=|d_{n,2}-d_{n,1}|+\cdots+|d_{n,m}-d_{n,m-1}|+|d_{n,m+1}-d_{n,m}|+\cdots+|d_{n,n}-d_{n,n-1}|$
$=d_{n,1}-d_{n,n},$
$\underline{R}_{2}$ $=|d_{2,2}-d_{2,1}|+\left(|d_{n-1,2}-d_{n-1,1}|+\cdots+|d_{n-1,n-1}-d_{n-1,n-2}|\right)$
$=(d_{2,1}-d_{2,2})+(d_{n-1,1}-d_{n-1,n-1}),$
$\vdots$
$\underline{R}_{m}$ $=(|d_{m,2}-d_{m,1}|+\cdots+|d_{m,m}-d_{m,m-1}|)+(|d_{m+1,2}-d_{m+1,1}|+\cdots+|d_{m+1,m+1}-d_{m+1,m}|)$
$=(d_{m,1}-d_{m,m})+(d_{m+1,1}-d_{m+1,m+1}).$
$C_{m,m}-B_{m,m}J_{m,m}+J_{m,m}A_{m,m}J_{m,m}$
$=\begin{bmatrix}d_{m+1,1}+d_{m,1}&d_{m,2}-d_{m,1}&&\cdots&d_{m,m}-d_{m,m-1}\\ d_{m+2,2}-d_{m+2,1}&d_{m+2,1}+d_{m-1,1}&&\cdots&d_{m-1,m-1}-d_{m-1,m-2}\\ \vdots&&\ddots&&\vdots\\ d_{n-1,m-1}-d_{n-1,m-2}&\cdots&&d_{n-1,1}+d_{2,1}&d_{2,2}-d_{2,1}\\ d_{n,m}-d_{n,m-1}&\cdots&&d_{n,2}-d_{n,1}&d_{n,1}+d_{1,1}\end{bmatrix}$
$-\begin{bmatrix}d_{m+1,2}-d_{m+1,1}&&\cdots&&d_{m+1,m+1}-d_{m+1,m}\\ \vdots&&\vdots&&\vdots\\ d_{n,m+1}-d_{n,m}&&\cdots&&d_{n,n}-d_{n,n-1}\end{bmatrix}$
Again for computational simplicity we take the Gershgorin radius from this block for the center $\underline{C}_{m+1}=d_{m+1,1}+d_{m,1},\underline{C}_{m+2}=d_{m+2,1}+d_{m-1,1},\cdots,\underline{C}_{n}=d_{n,1}+d_{1,1}$ are
$\underline{R}_{m+1}$ $=(|d_{m,2}-d_{m,1}|+\cdots+|d_{m,m}-d_{m,m-1}|)+(|d_{m+1,2}-d_{m+1,1}|+\cdots+|d_{m+1,m+1}-d_{m+1,m}|)$
$=(d_{m,1}-d_{m,m})+(d_{m+1,1}-d_{m+1,m+1}),$
$\vdots$
$\underline{R}_{n}$ $=|d_{n,2}-d_{n,1}|+\cdots+|d_{n,m}-d_{n,m-1}|+|d_{n,m+1}-d_{n,m}|+\cdots+|d_{n,n}-d_{n,n-1}|$
$=d_{n,1}-d_{n,n}.$
$A_{m,m}J_{m,m}-J_{m,m}C_{m,m}+J_{m,m}B_{m,m}J_{m,m}$
$=\begin{bmatrix}-(d_{n,m}-d_{n,m-1})&\cdots&&-(d_{n,2}-d_{n,1})&d_{1,1}-d_{n,1}\\ -(d_{n-1,m-1}-d_{n-1,m-2})&\cdots&&d_{2,1}-d_{n-1,1}&d_{2,2}-d_{2,1}\\ &&\iddots&&\vdots\\ -(d_{m+2,2}-d_{m+2,1})&d_{m-1,1}-d_{m+2,1}&&\cdots&d_{m-1,m-1}-d_{m-1,m-2}\\ d_{m,1}-d_{m+1,1}&d_{m,2}-d_{m,1}&&\cdots&d_{m,m}-d_{m,m-1}\end{bmatrix}$
$+\begin{bmatrix}d_{n,m+1}-d_{n,m}&&\cdots&&d_{n,n}-d_{n,n-1}\\ \vdots&&\vdots&&\vdots\\ d_{m+1,2}-d_{m+1,1}&&\cdots&&d_{m+1,m+1}-d_{m+1,m}\end{bmatrix}$
As this is anti-diagonal block so the contribution for the Gershgorin radius from this block is:
$\bar{R}_{1}$ $=\left(|d_{n,2}-d_{n,1}|+\cdots+|d_{n,m}-d_{n,m-1}|+|d_{n,m+1}-d_{n,m}|+\cdots+|d_{n,n}-d_{n,n-1}|\right)+|d_{1,1}-d_{n,1}|$
$=(d_{n,1}-d_{n,n})+|d_{1,1}-d_{n,1}|,$
$\bar{R}_{2}$ $=\left(|d_{n-1,2}-d_{n-1,1}|+\cdots+|d_{n-1,n-1}-d_{n-1,n-2}|\right)+|d_{2,1}-d_{n-1,1}|+|d_{2,2}-d_{2,1}|$
$=(d_{n-1,1}-d_{n-1,n-1})+|d_{2,1}-d_{n-1,1}|+(d_{2,1}-d_{2,2}),$
$\vdots$
$\bar{R}_{m}$ $=|d_{m,1}-d_{m+1,1}|+\left(|d_{m,2}-d_{m,1}|+\cdots+|d_{m,m}-d_{m,m-1}|\right)+\left(|d_{m+1,2}-d_{m+1,1}|+\cdots+|d_{m+1,m+1}-d_{m+1,m}|\right)$
$=(d_{m+1,1}-d_{m+1,m+1})+|d_{m,1}-d_{m+1,1}|+(d_{m,1}-d_{m,m}).$
$J_{m,m}A_{m,m}-C_{m,m}J_{m,m}-B_{m,m}$
$=\begin{bmatrix}d_{m,m}-d_{m,m-1}&\cdots&&d_{m,2}-d_{m,1}&d_{m,1}-d_{m+1,1}\\ d_{m-1,m-1}-d_{m-1,m-2}&\cdots&&d_{m-1,1}-d_{m+2,1}&-(d_{m+2,2}-d_{m+2,1})\\ &&\iddots&&\vdots\\ d_{2,2}-d_{2,1}&d_{2,1}-d_{n-1,1}&&\cdots&-(d_{n-1,m-1}-d_{n-1,m-2})\\ d_{1,1}-d_{n,1}&-(d_{n,2}-d_{n,1})&&\cdots&-(d_{n,m}-d_{n,m-1})\end{bmatrix}$
$-\begin{bmatrix}d_{m+1,m+1}-d_{m+1,m}&&\cdots&&d_{m+1,2}-d_{m+1,1}\\ \vdots&&\vdots&&\vdots\\ d_{n,n}-d_{n,n-1}&&\cdots&&d_{n,m+1}-d_{n,m}\end{bmatrix}$
Again this is anti-diagonal block so the contribution for the Gershgorin radius from this block is:
$\bar{R}_{m+1}$ $=(d_{m,1}-d_{m,m})+|d_{m,1}-d_{m+1,1}|+(d_{m+1,1}-d_{m+1,m+1}),$
$\vdots$
$\bar{R}_{n}$ $=|d_{1,1}-d_{n,1}|+(d_{n,1}-d_{n,n}).$
Hence when $n$ is even the Gershgorin radius of the matrix $X^{T}SX$ for center $C_{1}=\frac{d_{1,1}+d_{n,1}}{2},\cdots,C_{m}=\frac{d_{m,1}+d_{m+1,1}}{2},C_{m+1}=\frac{d_{m+1,1}+d_{m,1}}{2},\cdots,C_{n}=\frac{d_{n,1}+d_{1,1}}{2}$ are
$R_{1}$ $=(d_{n,1}-d_{n,n})+\frac{1}{2}|d_{1,1}-d_{n,1}|\leq\frac{3}{2}(d_{1,1}+d_{n,1}),$
$R_{2}$ $=(d_{n-1,1}-d_{n-1,n-1})+\frac{1}{2}|d_{2,1}-d_{n-1,1}|+(d_{2,1}-d_{2,2})\leq\frac{3}{2}(d_{2,1}+d_{n-1,1}),$
$\vdots$
$R_{m}$ $=(d_{m+1,1}-d_{m+1,m+1})+\frac{1}{2}|d_{m,1}-d_{m+1,1}|+(d_{m,1}-d_{m,m})\leq\frac{3}{2}(d_{m,1}+d_{m+1,1}),$
$R_{m+1}$ $=(d_{m,1}-d_{m,m})+\frac{1}{2}|d_{m,1}-d_{m+1,1}|+(d_{m+1,1}-d_{m+1,m+1})\leq\frac{3}{2}(d_{m,1}+d_{m+1,1}),$
$\vdots$
$R_{n}$ $=\frac{1}{2}|d_{1,1}-d_{n,1}|+(d_{n,1}-d_{n,n})\leq\frac{3}{2}(d_{1,1}+d_{n,1}).$
When $n$ is odd and $m:=\frac{n-1}{2}$ from relation 15 we have:
$X^{T}SX$
and
@@
and $B_{1,m}=[(d_{m+1,m+1}-d_{m+1,m})\,\cdots\,(d_{m+1,2}-d_{m+1,1})]$, $c=d_{m+1,1}$, $D_{m,1}=[(d_{m+2,2}-d_{m+2,1})\,\cdots\,(d_{n,m+1}-d_{n,m})]^{T}$. Hence block of $X^{T}SX$ are:
@@
For computational simplicity we take the Gershgorin radius from this block for the center $\underline{C}_{1}=d_{1,1}+d_{n,1},\underline{C}_{2}=d_{2,1}+d_{n-1,1},\cdots,\underline{C}_{m}=d_{m,1}+d_{m+2,1}$ are
$\underline{R}_{1}$ $=(|d_{n,2}-d_{n,1}|+\cdots+|d_{n,m}-d_{n,m-1}|)+(|d_{n,m+2}-d_{n,m+1}|+\cdots+|d_{n,n}-d_{n,n-1}|)$
@@
$-\begin{bmatrix}d_{m+2,3}-d_{m+2,2}&&\cdots&&d_{m+2,m+2}-d_{m+2,m+1}\\ \vdots&&\vdots&&\vdots\\ d_{n,m+2}-d_{n,m+1}&&\cdots&&d_{n,n}-d_{n,n-1}\end{bmatrix}$
Again for computational simplicity we take the Gershgorin radius from this block for the center $\underline{C}_{m+2}=d_{m+2,1}+d_{m,1},\underline{C}_{m+3}=d_{m+3,1}+d_{m-1,1},\cdots,\underline{C}_{n}=d_{n,1}+d_{1,1}$ are
@@
$+\begin{bmatrix}d_{n,m+2}-d_{n,m+1}&&\cdots&&d_{n,n}-d_{n,n-1}\\ \vdots&&\vdots&&\vdots\\ d_{m+2,3}-d_{m+2,2}&&\cdots&&d_{m+2,m+2}-d_{m+2,m+1}\end{bmatrix}$
, As this is anti-diagonal block so the contribution for the Gershgorin radius from this block is:
@@
Again this is anti-diagonal block so the contribution for the Gershgorin radius from this block are:
@@
$J_{m,m}D_{m,1}=[(d_{n,m+1}-d_{n,m})\,\cdots\,(d_{m+2,2}-d_{m+2,1})]^{T}$, $B_{1,m}J_{m,m}=[(d_{m+1,2}-d_{m+1,1})\,\cdots\,(d_{m+1,m+1}-d_{m+1,m})]$. Hence when $n$ is odd the Gershgorin radius of the matrix $X^{T}SX$ for center $C_{1}=\frac{d_{1,1}+d_{n,1}}{2},\cdots,C_{m}=\frac{d_{m,1}+d_{m+2,1}}{2},C_{m+1}=d_{m+1,1},C_{m+2}=\frac{d_{m+2,1}+d_{m,1}}{2},\cdots,C_{n}=\frac{d_{n,1}+d_{1,1}}{2}$ are
@@
Lemma 4.20.
Let $L=\begin{bmatrix}L_{1}&O\\ L_{2}&l\end{bmatrix}$ be a real lower triangular matrix of order $n+1$ and $L_{1}$ block is of order $n$. $B=\begin{bmatrix}B_{1}&O\\ O&0\end{bmatrix}$ be a real block matrix of order $n+1$ with block $B_{1}$ of order $n$ and $O$ be the zero matrix. If $(L_{1}-lI+\frac{i}{\sqrt{\sigma}}B_{1})$ is non-singular then the matrix $L+\frac{i}{\sqrt{\sigma}}B$ is similar to matrix $\bar{L}=\begin{bmatrix}L_{1}+\frac{i}{\sqrt{\sigma}}B_{1}&O\\ O&l\end{bmatrix}$.
Proof 4.21.
@@
$\begin{bmatrix}L_{1}+\frac{i}{\sqrt{\sigma}}D_{A}&0\\ L_{2}&l\end{bmatrix}\begin{bmatrix}X_{11}&X_{12}\\ X_{21}&X_{22}\end{bmatrix}$ $=\begin{bmatrix}X_{11}&X_{12}\\ X_{21}&X_{22}\end{bmatrix}\begin{bmatrix}L_{1}+\frac{i}{\sqrt{\sigma}}D_{A}&0\\ 0&l\end{bmatrix}$
$\begin{bmatrix}(L_{1}+\frac{i}{\sqrt{\sigma}}D_{A})X_{11}&(L_{1}+\frac{i}{\sqrt{\sigma}}D_{A})X_{12}\\ L_{2}X_{11}+lX_{21}&L_{2}X_{12}+lX_{22}\end{bmatrix}$ $=\begin{bmatrix}X_{11}(L_{1}+\frac{i}{\sqrt{\sigma}}D_{A})&X_{12}l\\ X_{21}(L_{1}+\frac{i}{\sqrt{\sigma}}D_{A})&X_{22}l\end{bmatrix}$
We choose $X_{11}=I,X_{12}=0,X_{22}=I$ then we have $X_{21}(L_{1}+\frac{i}{\sqrt{\sigma}}B_{1})=lX_{21}+L_{2}$. Therefore the system has solution if the matrix $(L_{1}-lI+\frac{i}{\sqrt{\sigma}}B_{1})$ is invertiable. Let $\underline{L}=(L_{1}-lI+\frac{i}{\sqrt{\sigma}}B_{1})^{-1}$ exist from our assumption. Therefor the matrix $L+\frac{i}{\sqrt{\sigma}}B$ is similar to $\bar{L}$ by the matrix $\begin{bmatrix}I&0\\ L_{2}+lL_{2}\underline{L}&1\end{bmatrix}$.
@@
Most of the practical cases $\sigma$ is very small therefore we can use lemma 4.16 and its corollaries 4.18 and 4.19 to show that $(L_{1}-lI+\frac{i}{\sqrt{\sigma}}B_{1})$ is invertiable when $B_{1}$ be the unit anti-diagonal matrix $J$. Again if $L_{1}-lI$ is toeplitz and $B_{1}=J$ then $|\det(L_{1}-lI+\frac{i}{\sqrt{\sigma}}J)|=|\det((L_{1}-lI)J+\frac{i}{\sqrt{\sigma}}I)|$. Now $(L_{1}-lI)J$ is symmetric so eigenvalues are all real hence $\det((L_{1}-lI)J+\frac{i}{\sqrt{\sigma}}I)\neq 0$ for any $\sigma$.
Before going to main convergence results for DNWR and NNWR algorithm for the control problem we want to introduce a different kind of mesh grading technique.

### v2

@@ 5 Auxiliary Results
$=\frac{1}{2}\begin{bmatrix}A_{m,m}+J_{m,m}C_{m,m}+J_{m,m}E_{m,m}J_{m,m}&\sqrt{2}J_{m,m}D_{m,1}&A_{m,m}J_{m,m}-J_{m,m}E_{m,m}+J_{m,m}C_{m,m}J_{m,m}\\ \sqrt{2}B_{1,m}&2cI_{1,1}&\sqrt{2}B_{1,m}J_{m,m}\\ J_{m,m}A_{m,m}-E_{m,m}J_{m,m}-C_{m,m}&\sqrt{2}D_{m,1}&J_{m,m}A_{m,m}J_{m,m}-C_{m,m}J_{m,m}+E_{m,m}\end{bmatrix}$,
where
@@
and $B_{1,m}=[(d_{m+1,m+1}-d_{m+1,m})\,\cdots\,(d_{m+1,2}-d_{m+1,1})]$, $c=d_{m+1,1}$, $D_{m,1}=[(d_{m+2,2}-d_{m+2,1})\,\cdots\,(d_{n,m+1}-d_{n,m})]^{T}$. Hence from the upper left block of $X^{T}SX$ i.e.
@@
the Gershgorin radius corresponding to the center $\bar{C}_{1}=d_{1,1}+d_{n,1},\bar{C}_{2}=d_{2,1}+d_{n-1,1},\cdots,\bar{C}_{m}=d_{m,1}+d_{m+2,1}$ are given by:
$\bar{R}_{1}$ $=(|d_{n,2}-d_{n,1}|+\cdots+|d_{n,m}-d_{n,m-1}|)+(|d_{n,m+2}-d_{n,m+1}|+\cdots+|d_{n,n}-d_{n,n-1}|)$
@@
$-\begin{bmatrix}d_{m+2,3}-d_{m+2,2}&&\cdots&&d_{m+2,m+2}-d_{m+2,m+1}\\ \vdots&&\vdots&&\vdots\\ d_{n,m+2}-d_{n,m+1}&&\cdots&&d_{n,n}-d_{n,n-1},\end{bmatrix}$
we take the Gershgorin radius corresponding to the center $\underline{C}_{m+2}=d_{m+2,1}+d_{m,1},\underline{C}_{m+3}=d_{m+3,1}+d_{m-1,1},\cdots,\underline{C}_{n}=d_{n,1}+d_{1,1}$ as
@@
$+\begin{bmatrix}d_{n,m+2}-d_{n,m+1}&&\cdots&&d_{n,n}-d_{n,n-1}\\ \vdots&&\vdots&&\vdots\\ d_{m+2,3}-d_{m+2,2}&&\cdots&&d_{m+2,m+2}-d_{m+2,m+1}\end{bmatrix}$,
the contribution to the Gershgorin radius are:
@@
the contribution to the Gershgorin radius are:
@@
Also we have $J_{m,m}D_{m,1}=[(d_{n,m+1}-d_{n,m})\,\cdots\,(d_{m+2,2}-d_{m+2,1})]^{T}$, $B_{1,m}J_{m,m}=[(d_{m+1,2}-d_{m+1,1})\,\cdots\,(d_{m+1,m+1}-d_{m+1,m})]$. Hence the Gershgorin radius of the matrix $X^{T}SX$ corresponding to the center $C_{1}=\frac{d_{1,1}+d_{n,1}}{2},\cdots,C_{m}=\frac{d_{m,1}+d_{m+2,1}}{2},C_{m+1}=d_{m+1,1},C_{m+2}=\frac{d_{m+2,1}+d_{m,1}}{2},\cdots,C_{n}=\frac{d_{n,1}+d_{1,1}}{2}$ are given by:
@@
Lemma 10.
Let $S=\begin{bmatrix}S_{1}&O\\ S_{2}&s\end{bmatrix}$ be a real lower triangular matrix of order $n+1$ and $S_{1}$ block is of order $n$. $B=\begin{bmatrix}B_{1}&O\\ O&0\end{bmatrix}$ be a real block matrix of order $n+1$ with block $B_{1}$ of order $n$ and $O$ be the zero matrix. If $(S_{1}-sI+i\beta B_{1})$ is non-singular for positive constant $\beta$ then the matrix $S+i\beta B$ is similar to matrix $\bar{S}=\begin{bmatrix}S_{1}+i\beta B_{1}&O\\ O&s\end{bmatrix}$.
Proof.
@@
$\begin{bmatrix}S_{1}+i\beta B_{1}&O\\ S_{2}&s\end{bmatrix}\begin{bmatrix}X_{11}&X_{12}\\ X_{21}&X_{22}\end{bmatrix}$ $=\begin{bmatrix}X_{11}&X_{12}\\ X_{21}&X_{22}\end{bmatrix}\begin{bmatrix}S_{1}+i\beta B_{1}&O\\ O&s\end{bmatrix}$
$\begin{bmatrix}(S_{1}+i\beta B_{1})X_{11}&(S_{1}+i\beta B_{1})X_{12}\\ S_{2}X_{11}+sX_{21}&S_{2}X_{12}+sX_{22}\end{bmatrix}$ $=\begin{bmatrix}X_{11}(S_{1}+i\beta B_{1})&X_{12}l\\ X_{21}(S_{1}+i\beta B_{1})&X_{22}l\end{bmatrix}$
We choose $X_{11}=I,X_{12}=O,X_{22}=1$ then we have $X_{21}(S_{1}+i\beta B_{1})=sX_{21}+S_{2}$. Therefore the system has solution if the matrix $(S_{1}-sI+i\beta B_{1})$ is invertiable. Hence the matrix $S+i\beta B$ is similar to $\bar{S}$ by the matrix $\begin{bmatrix}I&O\\ S_{2}(S_{1}-sI+i\beta B_{1})^{-1}&1\end{bmatrix}$. ∎
@@
If $S_{1}$ be toeplitz and $B_{1}$ be the exchange matrix $J$ then $|\det(S_{1}-sI+i\beta J)|=|\det((S_{1}-sI)J+i\beta I)|$. Now $(S_{1}-sI)J$ is symmetric so eigenvalues are all real hence $\det((S_{1}-sI)J+i\beta I)\neq 0$ for any $\sigma$.
Before going to main convergence results for DNWR and NNWR algorithm for the control problem we want to introduce a different kind of mesh grading technique, which is efficient and also helpful for the analysis.